# Afford2Act: Affordance-Guided Automatic Keypoint Selection for Generalizable and Lightweight Robotic Manipulation

### v1

@@ V Conclusions and Future Work
We introduced Afford2Act, a three-stage imitation learning pipeline that leverages semantically meaningful keypoints derived from affordance text prompts. Through six real-world experiments, we demonstrated that Afford2Act is not only data-efficient and performant, but also generalizes effectively to unseen objects, categories, backgrounds, and distractors. While our work represents the first attempt at semantically meaningful keypoint-based imitation learning, the selected keypoints remain fixed during inference and cannot be adapted dynamically to evolving contexts. As discussed in Sec. IV-G, integrating task planning with keypoint resampling is a promising future direction, potentially enabling flexible long-horizon policy execution.

### v2

@@ V Conclusions and Future Work
We present Afford2Act, an imitation learning framework that uses language-conditioned affordance cues to distill a compact set of semantically grounded 2D keypoints for manipulation from a single external RGB view. Across six real-robot tasks, Afford2Act demonstrates data-efficient learning and robust performance under real-world variation, including unseen instances, novel categories with shared functional parts, background changes, and distractors. This indicates that affordance-grounded keypoints can serve as an effective lightweight state representation for manipulation. A promising direction is to allow the keypoint representation to update online as the scene evolves. As discussed in Sec. IV-G, integrating lightweight task planning with confidence-aware keypoint resampling could improve adaptability during long-horizon execution, particularly under changing visibility or contact.